Benzene has a heat of vaporization of 30.72 kJ/mol and a normal boiling point of 80.1 ∘C. At what temperature does benzene boil when the external pressure is 410 torr?


1
Expert's answer
2021-09-01T03:11:23-0400
  • The external pressure of the benzene is:

Pb = 410 torr = 54662.20 Pb

  • The molar enthalpy of vaporization of benzene is:

ΔHa =30.72 Kj/mol = 30720 J/mol

  • The normal boiling temperature of the benzene is:

Ta = 80.10oC = 353.10K

The temperature of benzene to boil at 410 torr is Tb


atmospheric pressure is 101325 Pa


The benzene boil at normal atmospheric pressure where boiling temperature is 80.10


The expression by using the Clausius-Clapeyron equation to evaluate the temperature of benzene to boil at pressure of 410 torr is


  • ln( 101325 /54662.20 ) = (30720 / 8.314). (1/Tb - 1/ 353.10 )
  • (0.61716×8.314 / 30720 ) = 1/ Tb - 1/353.10 )
  • 1/Tb = 1.670 x 10-4 + 1/353.10
  • 1/Tb = 1.670 x 10-4 + 0.002832
  • 1/Tb = 0.002999
  • Tb = 1/0.002999
  • Tb = 333.44 K
  • Tb = 60.44oC



Thus the boiling temperature of benzene at pressure of 410 torr is 60.44oC